find all complex number Z such that |Z|=1 and |z/z-+z-/z-|=1


  1. find all complex number Z such that |Z|=1 and |z/z-+z-/z-|=1

Grade:Upto college level

1 Answers

Priyansh Bajaj AskiitiansExpert-IITD
13 years ago

Dear Sachin,


|(z/z-) + (z-/z-)| =1  means |(z/z-) + 1| =1; which further means (z/z-) =0 or z =0

But, it is given that |z| =1. So, z can't be zero.

Therefore, you've typed the wrong question. Please check the question and post it again(if you've any doubt).


Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

All the best Sachin!!!


Regards,

Askiitians Experts

Priyansh Bajaj